### v1

# Subsystem fidelity in two-dimensional conformal field theories

Abstract
We investigate the short-interval expansion of the subsystem fidelity in two-dimensional conformal field theories (2D CFTs) using the operator product expansion (OPE) of twist operators. We obtain universal contributions from general quasiprimary operators valid for arbitrary 2D CFTs, along with specific results in free massless boson and fermion theories. The analytical predictions demonstrate excellent agreement with established analytical results in field theories and numerical calculations in integrable models. Furthermore, we extend the method to holographic CFTs, where subsystem fidelity serves to analyze the distinguishability of black hole microstates through the AdS/CFT correspondence. This work establishes a unified framework for quantifying quantum state distinguishability across various 2D CFTs, bridging quantum information techniques with applications in quantum gravity.
Center for Joint Quantum Studies and Department of Physics, School of Science, Tianjin University, 135 Yaguan Road, Tianjin 300350, China
1 Introduction
The quantitative distinguishability between quantum states plays a fundamental role across multiple domains of theoretical physics. Quantitative measures of distinguishability not only serve as operational tools but also bridge fundamental concepts between quantum information, many-body physics, and quantum gravity. In quantum information theory, it provides a crucial metric for evaluating the precision and reliability of quantum states in various operational tasks, such as quantum metrology and quantum computing, where optimal discrimination between states directly impacts protocol performance [1, 2]. Moreover, in the study of quantum thermalization, state distinguishability offers a powerful diagnostic tool to characterize and compare thermalization processes in chaotic versus integrable quantum systems, thereby offering insights into the emergence of statistical mechanics from unitary quantum dynamics [3, 4, 5, 6]. Furthermore, this concept has been rigorously formulated within quantum field theory [7, 8, 9, 10], where it helps quantify the distinguishability of eigenstates and states under unitary evolution, with significant implications for understanding the structure of the Hilbert space. In the context of the AdS/CFT correspondence [11, 12, 13], state distinguishability is intimately related to the problem of bulk reconstruction, determining how boundary quantum states encode information about the geometry and fields in the anti-de Sitter bulk spacetime [14, 15, 16, 17].
In this paper, we focus on calculating subsystem fidelity in two-dimensional conformal field theory (2D CFT). Fidelity measures the similarity between two quantum states. For two density matrices $\rho$ and $\sigma$, it is defined as [1, 2]
$F(\rho,\sigma)=\textrm{tr}\sqrt{\sqrt{\rho},\sigma\sqrt{\rho}}.$ (1.1)
Direct evaluation of this expression with a large Hilbert space is often challenging due to the presence of square roots. Inspired by the replica trick used in entanglement entropy calculations [18, 19], several replica-based approaches have been developed to compute fidelity. In [7], the following replica expression was used
$F(\rho,\sigma)=\lim_{p\to 1/2}\textrm{tr}[(\rho^{\frac{1-p}{2p}}\sigma\rho^{\frac{1-p}{2p}})^{p}].$ (1.2)
A simplified two-parameter replica trick was later used in [16, 17]
$F(\rho,\sigma)=\lim_{m\to 1/2,,n\to 1/2}\textrm{tr}[(\rho^{m}\sigma\rho^{m})^{n}].$ (1.3)
A more computationally efficient formula of fidelity was suggested in [20]
$F(\rho,\sigma)=\textrm{tr}\sqrt{\rho\sigma}.$ (1.4)
More recently, a refined replica trick was introduced and applied in the context of AdS/CFT correspondence in [21]
$F(\rho,\sigma)=\lim_{p\to 1/2}\textrm{tr}[(\rho\sigma)^{p}].$ (1.5)
In this work, we adopt this last replica trick to evaluate subsystem fidelity in 2D CFTs.
We study subsystem fidelity in general 2D CFTs using twist operators [19, 22, 23] and their operator product expansion (OPE) [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 9, 10]. By evaluating the short-interval expansion of subsystem fidelity via the OPE of twist operators, we show that it receives contributions from various quasiprimary operators ordered by their scaling dimensions. We derive universal contributions from general quasiprimary operators in arbitrary 2D CFTs, as well as specific contributions from particular operators in free boson and free fermion theories. Our results in these solvable models agree with existing analytical and numerical calculations. Furthermore, we extend the method to 2D holographic CFTs, where through the AdS3/CFT2 correspondence [35], we analyze the perturbative distinguishability of black hole microstates.
The paper is organized as follows. In Section 2, we outline the general method for computing subsystem fidelity using the short-interval expansion of twist operator OPEs, including both universal and model-specific contributions. In Sections 3 and 4, we validate the approach by calculating subsystem fidelity between low-lying eigenstates in the 2D massless free boson and fermion theories, respectively. Section 5 applies the framework to holographic CFTs, exploring the distinguishability of black hole microstates. We conclude in Section 6 with a summary and discussion. Additional technical details on the analytical continuations used in the fidelity calculations are provided in Appendix A.
2 General method
In this section, we outline the general method for calculating the short-interval expansion of subsystem fidelity using the OPE of twist operators in 2D CFTs. We begin by reviewing relevant basic concepts, then apply the replica trick to derive subsystem fidelity expressed in terms of contributions from various quasiprimary operators in the replicated CFT. We discuss several universal contributions to subsystem fidelity in general 2D CFTs and two specific types of particular contributions in free massless boson and fermion theories.
2.1 Relevant basics
In this subsection, we review essential background concepts of 2D CFTs. For more details, see [36, 37, 38].
From global conformal symmetry, all independent operators in a general 2D CFT can be classified as primary operators and their descendants, while from the global SL(2,C)(2,C) conformal symmetry, all operators can be classified as quasiprimary operators and their derivatives. Under a general local conformal transformation $z\to f(z)$, a primary operator $\phi$ with conformal weights $(h_{\phi},\bar{h}_{\phi})$ transforms as
$\phi(z,\bar{z})=f^{\prime}(z)^{h_{\phi}}\bar{f}^{\prime}(\bar{z})^{\bar{h}_{\phi}}\phi(f(z),\bar{f}(\bar{z})),$ (2.1)
while a quasiprimary operator $\phi$ transforms as
$\phi(z,\bar{z})=f^{\prime}(z)^{h_{\phi}}\bar{f}^{\prime}(\bar{z})^{\bar{h}_{\phi}}\phi(f(z),\bar{f}(\bar{z}))+\cdots,$ (2.2)
where $\cdots$ denotes terms containing the Schwarz derivative
$s(z)=\frac{f^{\prime\prime\prime}(z)}{f^{\prime}(z)}-\frac{3}{2}\Big(\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\Big)^{2},$ (2.3)
and/or its derivatives. From the conformal weights $(h_{\phi},\bar{h}_{\phi})$, one obtains the scaling dimension $\Delta_{\phi}=h_{\phi}+\bar{h}_{\phi}$ and spin $s_{\phi}=h_{\phi}-\bar{h}_{\phi}$. The operator $\phi$ is bosonic when $s_{\phi}$ is an integer and fermionic when $s_{\phi}$ is half-integer. In both cases, $\mathrm{i}^{8s_{\phi}}=1$.
The correlation function for two quasiprimary operators on a complex plane $C$ is
$\langle\phi(z_{1},\bar{z}_{1})\psi(z_{2},\bar{z}_{2})\rangle_{C}=\frac{\alpha_{\phi}\delta_{\phi\psi}}{z_{12}^{2h_{\phi}}\bar{z}_{12}^{2\bar{h}_{\phi}}},$ (2.4)
where $\alpha_{\phi}$ relates to the normalization of operator $\phi$, all quasiprimary operators have been orthogonalized, and we use $z_{12}\equiv z_{1}-z_{2}$ and $\bar{z}_{12}\equiv\bar{z}_{1}-\bar{z}_{2}$. The correlation function of three quasiprimary operators on a plane is
$\langle\phi(z_{1},\bar{z}_{1})\psi(z_{2},\bar{z}_{2})\chi(z_{3},\bar{z}_{3})\rangle_{C}=\frac{C_{\phi\psi\chi}}{z_{12}^{h_{\phi}+h_{\psi}-h_{\chi}}z_{13}^{h_{\phi}+h_{\chi}-h_{\psi}}z_{23}^{h_{\psi}+h_{\chi}-h_{\phi}}\bar{z}_{12}^{\bar{h}_{\phi}+\bar{h}_{\psi}-\bar{h}_{\chi}}\bar{z}_{13}^{\bar{h}_{\phi}+\bar{h}_{\chi}-\bar{h}_{\psi}}\bar{z}_{23}^{\bar{h}_{\psi}+\bar{h}_{\chi}-\bar{h}_{\phi}}},$ (2.5)
where $C_{\phi\psi\chi}$ is the structure constant.
Typical quasiprimary operators include the stress tensor $T$ with conformal weights $(2,0)$ and normalization $\alpha_{T}=\frac{c}{2}$, where $c$ is the central charge of the 2D CFT, and the operator
$\mathcal{A}=(TT)-\frac{3}{10}\partial^{2}T,$ (2.6)
with conformal weights $(4,0)$ and normalization
$\alpha_{\mathcal{A}}=\frac{c(5c+22)}{10}.$ (2.7)
Similar considerations apply to the anti-holomorphic stress tensor $\bar{T}$ with conformal weights $(0,2)$ and the operator $\bar{\mathcal{A}}=(\bar{T}\bar{T})-\frac{3}{10}\bar{\partial}^{2}\bar{T}$ with conformal weights $(0,4)$.
A pure state $|\phi\rangle$ in a 2D CFT on a cylinder with coordinate $w$ and spatial period $w\sim w+L$ corresponds to inserting an operator $\phi(z,\bar{z})$ at the origin on a plane with coordinate $z=\mathrm{e}^{\frac{2\pi\mathrm{i}w}{L}}$
$|\phi\rangle=\phi(0,0)|G\rangle,$ (2.8)
with $|G\rangle$ denoting the ground state. When $\phi$ is a primary operator, we have the expectation values
$\langle T\rangle_{\phi}=\frac{\pi^{2}(c-24h_{\phi})}{6L^{2}},~~\langle\mathcal{A}\rangle_{\phi}=\frac{\pi^{4}[c(5c+22)-240(c+2)h_{\phi}+2880h_{\phi}^{2}]}{180L^{4}}.$ (2.9)
It is similar for $\langle\bar{T}\rangle_{\phi}$ and $\langle\bar{\mathcal{A}}\rangle_{\phi}$.
We also consider the thermal state
$\rho_{\beta}=\frac{\mathrm{e}^{-\beta H}}{Z(\beta)},$ (2.10)
where $\beta$ is the inverse temperature, $H$ is the Hamiltonian, and $Z(\beta)=\textrm{tr}(\mathrm{e}^{-\beta H})$ is the partition function. Useful thermal state expectation values are
$\langle T\rangle_{\beta}=-\frac{\pi^{2}c}{6\beta^{2}},~~\langle\mathcal{A}\rangle_{\beta}=\frac{\pi^{4}c(5c+22)}{180\beta^{4}}.$ (2.11)
It is the same for $\langle\bar{T}\rangle_{\beta}$ and $\langle\bar{\mathcal{A}}\rangle_{\beta}$.
2.2 OPE of twist operators
We consider one interval $A=[0,\ell]$ on a circle of length $L$ in two general states $\rho$ and $\sigma$ in a general 2D CFT. The subsystem fidelity can be calculated as [20]
$F(\rho_{A},\sigma_{A})=\textrm{tr}_{A}\sqrt{\rho_{A}\sigma_{A}}.$ (2.12)
We define the “generalized fidelity” as
$F_{p}(\rho_{A},\sigma_{A})=\textrm{tr}_{A}[(\rho_{A}\sigma_{A})^{p}],$ (2.13)
and obtain fidelity from the limit [21]
$F(\rho_{A},\sigma_{A})=\lim_{p\to 1/2}F_{p}(\rho_{A},\sigma_{A}).$ (2.14)
The replica approach proceeds as follows: we first evaluate the generalized fidelity $F_{p}(\rho_{A},\sigma_{A})$ for general positive integer values of $p=1,2,\cdots$, corresponding to positive even integer values of $n=2p=2,4,\cdots$, and then apply analytical continuation $p\to 1/2$ (i.e., $n\to 1$) to obtain the fidelity.
In this paper, we only consider translationally invariant states. Using twist operators $\mathcal{T}$ and $\tilde{\mathcal{T}}$ [19, 22, 23] and their OPE [24, 25, 26, 27, 28, 29, 30, 31, 32, 34, 33, 9, 10], we obtain the short interval expansion
$\langle\mathcal{T}(\ell,\ell)\tilde{\mathcal{T}}(0,0)\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{2(h_{n}+\bar{h}_{n})}\Big(1+\sum_{K}\ell^{\Delta_{K}}d_{K}\langle\Phi_{K}\rangle_{\rho\otimes\sigma\otimes\rho\otimes\sigma\cdots}\Big).$ (2.15)
The twist operators $\mathcal{T}$ and $\tilde{\mathcal{T}}$ are primary operators in the replicated $n$-fold CFT, which we denote as CFTn, and their conformal weights are [19, 23]
$h_{n}=\bar{h}_{n}=\frac{c(n^{2}-1)}{24n},$ (2.16)
where $c$ is the central charge of the original single copy CFT. The factor $c_{n}$ relates to the normalization of twist operators, satisfying $\lim_{n\to 1}c_{n}=1$, and $\epsilon$ is the UV cutoff. Due to translational invariance, we only need to consider CFTn non-identity quasiprimary operators $\Phi_{K}$ that are direct products of quasiprimary operators in different replicas
$\Phi_{K}=\mathcal{X}_{1}^{j_{1}}\mathcal{X}_{2}^{j_{2}}\cdots\mathcal{X}_{k}^{j_{k}}.$ (2.17)
The set $\{\mathcal{X}_{1},\cdots,\mathcal{X}_{k}\}$ consists of non-identity quasiprimary operators in the original single copy of the CFT. Here $0\leq j_{1},j_{2},\cdots,j_{k}\leq n-1$ are replica indices taking appropriate values to avoid undercounting or overcounting of CFTn quasiprimary operators. Also because of translational invariance, terms involving the derivatives of the CFTn quasiprimary operators $\Phi_{K}$ do not appear in (2.15).
We obtain the expansion of generalized fidelity
$F_{p}(\rho_{A},\sigma_{A})=c_{n}\Big(\frac{\ell}{\epsilon}\Big)^{2(h_{n}+\bar{h}_{n})}\bigg(1+\sum_{k=1}^{n}\sum_{\{\mathcal{X}_{1},\cdots,\mathcal{X}_{k}\}}\ell^{\Delta_{\mathcal{X}_{1}}+\cdots+\Delta_{\mathcal{X}_{k}}}\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(p)}\bigg),$ (2.18)
where we define
$\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(p)}\equiv\sum_{j_{1},\cdots,j_{k}}d_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{j_{1}\cdots j_{k}}\langle\mathcal{X}_{1}\rangle_{j_{1}}\cdots\langle\mathcal{X}_{k}\rangle_{j_{k}},$ (2.19)
and the expectation value
$\langle\mathcal{X}\rangle_{j}\equiv\left\{\begin{array}[]{cl}\langle\mathcal{X}\rangle_{\rho}&j\rm{~is~even}\\ \langle\mathcal{X}\rangle_{\sigma}&j\rm{~is~odd}\end{array}\right.\!\!\!$. (2.20)
Note that $\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(p)}$ depends on states $\rho,\sigma$, which we omit for conciseness. For $k=1$, there is always $d_{\mathcal{X}}=O(n-1)$ [34], so $\mathcal{F}_{\mathcal{X}}^{(1/2)}=0$. We obtain the short interval expansion of subsystem fidelity
$F(\rho_{A},\sigma_{A})=1+\sum_{k=2}^{+\infty}\sum_{\mathcal{X}_{1},\cdots,\mathcal{X}_{k}}\ell^{\Delta_{\mathcal{X}_{1}}+\cdots+\Delta_{\mathcal{X}_{k}}}\mathcal{F}_{\mathcal{X}_{1}\cdots\mathcal{X}_{k}}^{(1/2)}.$ (2.21)
This expansion provides a systematic way to compute subsystem fidelity order by order in the interval length $\ell$, with contributions from various combinations of quasiprimary operators.
2.3 Universal contributions from $\mathcal{X}\mathcal{X}$
In this subsection, we derive the universal contributions to subsystem fidelity from pairs of identical quasiprimary operators $\{\mathcal{X},\mathcal{X}\}$. We have a quasiprimary operator $\mathcal{X}$ with normalization $\alpha_{\mathcal{X}}$, scaling dimension $\Delta_{\mathcal{X}}$, and spin $s_{\mathcal{X}}$. Note that $\mathcal{X}$ needs not be primary.
The OPE coefficient for the CFTn operator $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}$ with $0\leq j_{1}<j_{2}\leq n-1$ is given by [25, 39, 40]
$d_{\mathcal{X}\mathcal{X}}^{j_{1}j_{2}}=\frac{\mathrm{i}^{2s_{\mathcal{X}}}}{(2n)^{2\Delta_{\mathcal{X}}}\alpha_{\mathcal{X}}}\frac{1}{\big|\sin\frac{\pi j_{12}}{n}\big|^{2\Delta_{\mathcal{X}}}}+O(n-1),$ (2.22)
with shorthand $j_{12}\equiv j_{1}-j_{2}$. The O(n1)O(n-1) part exists if $\mathcal{X}$ is quasiprimary but not primary. Since we take the $n\to 1$ limit for the final result, the O(n1)O(n-1) part does not contribute to fidelity. The contribution from $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}$ is the $p\to\frac{1}{2}$ limit (i.e., $n\to 1$ limit) of
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}}=\sum_{0\leq j_{1}<j_{2}\leq n-1}d_{\mathcal{X}\mathcal{X}}^{j_{1}j_{2}}\langle\mathcal{X}\rangle_{j_{1}}\langle\mathcal{X}\rangle_{j_{2}},$ (2.23)
where $j_{12}\equiv j_{1}-j_{2}$.
To evaluate this sum, we separate the replica indices into even and odd values. Denoting even integers $a=0,2,\cdots,2p-2$ and odd integers $b=1,3,\cdots,2p-1$, we get
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle\rho^{2}+\langle\mathcal{X}\rangle_{\sigma}^{2})\sum_{a_{1}<a_{2}}d_{\mathcal{X}\mathcal{X}}^{a_{1}a_{2}}+\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a,b}d_{\mathcal{X}\mathcal{X}}^{ab}.$ (2.24)
Noting the identity
$2\sum_{a_{1}<a_{2}}d_{\mathcal{X}\mathcal{X}}^{a_{1}a_{2}}+\sum_{a,b}d_{\mathcal{X}\mathcal{X}}^{ab}=O(n-1),$ (2.25)
we can simplify the expression to
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}}=\frac{\mathrm{i}^{2s\mathcal{X}}}{2^{2\Delta_{\mathcal{X}}}\alpha_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}G_{1}(p,\Delta_{\mathcal{X}})+O(n-1),$ (2.26)
where the function $G_{1}(p,\Delta_{\mathcal{X}})$ is defined in (A.1) and encapsulates the sum over even replica indices. Using the analytical continuation (A.5), we obtain the universal contributions from $\mathcal{X}\mathcal{X}$ to subsystem fidelity
$\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}=-\frac{\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})}{2^{2\Delta_{\mathcal{X}}+3}\sqrt{\pi}\Gamma(\Delta_{\mathcal{X}}+1)}\frac{\mathrm{i}^{2s_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}}{\alpha_{\mathcal{X}}},$ (2.27)
as reported in [41].
For $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}}$ to be nonvanishing, the operator $\mathcal{X}$ must be bosonic, meaning the spin $s_{\mathcal{X}}$ is an integer, satisfying $\mathrm{i}^{4s_{\mathcal{X}}}=1$. This result shows that the leading contribution from pairs of identical operators depends quadratically on the difference of their expectation values in the two states.
2.4 Universal contributions from $\mathcal{X}\mathcal{X}\mathcal{X}$
We now consider contributions from triple products of identical quasiprimary operators $\mathcal{X}\mathcal{X}\mathcal{X}$. The OPE coefficient for the CFTn quasiprimary operator $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}\mathcal{X}_{j_{3}}$ with $0\leq j_{1}<j_{2}<j_{3}\leq n-1$ is [25, 39, 40]
$d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{j_{1}j_{2}j_{3}}=\frac{\mathrm{i}^{s_{\mathcal{X}}}C_{\mathcal{X}\mathcal{X}\mathcal{X}}}{(2n)^{3\Delta_{\mathcal{X}}}\alpha_{\mathcal{X}}^{3}}\frac{1}{\big|\sin\frac{\pi j_{12}}{n}\sin\frac{\pi j_{13}}{n}\sin\frac{\pi j_{23}}{n}\big|^{\Delta_{\mathcal{X}}}}+O(n-1).$ (2.28)
The structure constant $C_{\mathcal{X}\mathcal{X}\mathcal{X}}$ vanishes unless the spin $s_{\mathcal{X}}$ is an even integer [38]. The contributions from $\mathcal{X}_{j_{1}}\mathcal{X}_{j_{2}}\mathcal{X}_{j_{3}}$ to fidelity correspond to the $n\to 1$ limit of
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=\sum_{0\leq j_{1}<j_{2}<j_{3}\leq n-1}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{j_{1}j_{2}j_{3}}\langle\mathcal{X}\rangle_{j_{1}}\langle\mathcal{X}\rangle_{j_{2}}\langle\mathcal{X}\rangle_{j_{3}}.$ (2.29)
Separating the sum into contributions from different combinations of even and odd replica indices, we obtain
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=(\langle\mathcal{X}\rangle\rho^{3}+\langle\mathcal{X}\rangle_{\sigma}^{3})\sum_{a_{1}<a_{2}<a_{3}}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}a_{3}}+(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})\langle\mathcal{X}\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}\sum_{a_{1}<a_{2},b}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}b}.$ (2.30)
Using the identity
$\sum_{a_{1}<a_{2}<a_{3}}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}a_{3}}+\sum_{a_{1}<a_{2},b}d_{\mathcal{X}\mathcal{X}\mathcal{X}}^{a_{1}a_{2}b}=O(n-1),$ (2.31)
we simplify to
$\mathcal{F}^{(p)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=\frac{\mathrm{i}^{s\mathcal{X}}C_{\mathcal{X}\mathcal{X}\mathcal{X}}}{2^{3\Delta_{\mathcal{X}}}\alpha_{\mathcal{X}}^{3}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})G_{2}(p,\Delta_{\mathcal{X}})+O(n-1),$ (2.32)
where the function $G_{2}(p,\Delta_{\mathcal{X}})$ is defined in (A.6). Using the analytical continuation (A.8), we obtain the universal contributions from $\mathcal{X}\mathcal{X}\mathcal{X}$ to subsystem fidelity
$\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}=\frac{\Gamma(\Delta\mathcal{X}+\frac{1}{2})^{2}}{2^{2\Delta_{\mathcal{X}}+4}\pi\Gamma(\frac{\Delta_{\mathcal{X}}}{2}+1)\Gamma(\frac{3\Delta_{\mathcal{X}}}{2}+1)}\frac{\mathrm{i}^{s_{\mathcal{X}}}C_{\mathcal{X}\mathcal{X}\mathcal{X}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}(\langle\mathcal{X}\rangle_{\rho}+\langle\mathcal{X}\rangle_{\sigma})}{\alpha_{\mathcal{X}}^{3}}.$ (2.33)
For $\mathcal{F}^{(1/2)}_{\mathcal{X}\mathcal{X}\mathcal{X}}$ to be nonvanishing, $s_{\mathcal{X}}$ must be an even integer, satisfying $\mathrm{i}^{2s_{\mathcal{X}}}=1$. This cubic contribution provides the sub-leading order correction to the fidelity beyond the quadratic terms.
2.5 Universal contributions from $T\mathcal{X}\mathcal{X}$
We now examine contributions involving the stress tensor $T$ combined with two identical quasiprimary operators $\mathcal{X}$. The OPE coefficient for the operator $T_{j_{1}}\mathcal{X}_{j_{2}}\mathcal{X}_{j_{3}}$ with $j_{1}\neq j_{2}$, $j_{1}\neq j_{3}$, $j_{2}<j_{3}$ is [39, 40]
$d_{T\mathcal{X}\mathcal{X}}^{j_{1}j_{2}j_{3}}=-\frac{2\mathrm{i}^{2s_{\mathcal{X}}}h_{\mathcal{X}}}{(2n)^{2\Delta_{\mathcal{X}}+2}c\alpha_{\mathcal{X}}}\frac{1}{\big(\sin\frac{\pi j_{12}}{n}\sin\frac{\pi j_{13}}{n}\big)^{2}\big|\sin\frac{\pi j_{23}}{n}\big|^{2\Delta_{\mathcal{X}}-2}}+O(n-1).$ (2.34)
After summing over replica indices and separating contributions, we obtain
$\mathcal{F}_{T\mathcal{X}\mathcal{X}}^{(p)}=-\frac{\mathrm{i}^{2s_{\mathcal{X}}}h_{\mathcal{X}}}{2^{2\Delta_{\mathcal{X}}+1}c\alpha_{\mathcal{X}}}\big[(\langle T\rangle_{\rho}\langle\mathcal{X}\rangle_{\rho}^{2}+\langle T\rangle_{\sigma}\langle\mathcal{X}\rangle_{\sigma}^{2})G_{3}(p,\Delta_{\mathcal{X}})+(\langle T\rangle_{\rho}\langle\mathcal{X}\rangle_{\sigma}^{2}+\langle T\rangle_{\sigma}\langle\mathcal{X}\rangle_{\rho}^{2})G_{4}(p,\Delta_{\mathcal{X}})$
[…]
where $G_{3}(p,\Delta_{\mathcal{X}})$, $G_{4}(p,\Delta_{\mathcal{X}})$, and $G_{5}(p,\Delta_{\mathcal{X}})$ are defined in (A.9), (A.12), and (A.15), respectively. Using the analytical continuations (A.11), (A.14), and (A.17), we obtain the universal contributions of $T\mathcal{X}\mathcal{X}$ to subsystem fidelity
$\mathcal{F}_{T\mathcal{X}\mathcal{X}}^{(1/2)}=\frac{h_{\mathcal{X}}\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})}{2^{2\Delta_{\mathcal{X}}+5}\sqrt{\pi}c\Gamma(\Delta_{\mathcal{X}}+2)}\frac{\mathrm{i}^{2s_{\mathcal{X}}}}{\alpha_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})\big[(\langle T\rangle_{\rho}+\langle T\rangle_{\sigma})(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})$
[…]
Similarly, we obtain contributions of $\bar{T}\mathcal{X}\mathcal{X}$ to subsystem fidelity
$\mathcal{F}_{\bar{T}\mathcal{X}\mathcal{X}}^{(1/2)}=\frac{\bar{h}_{\mathcal{X}}\Gamma(\Delta_{\mathcal{X}}+\frac{1}{2})}{2^{2\Delta_{\mathcal{X}}+5}\sqrt{\pi}c\Gamma(\Delta_{\mathcal{X}}+2)}\frac{\mathrm{i}^{2s_{\mathcal{X}}}}{\alpha_{\mathcal{X}}}(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})\big[(\langle\bar{T}\rangle_{\rho}+\langle\bar{T}\rangle_{\sigma})(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})$ (2.37)
[…]
These mixed contributions involving the stress tensor provide important corrections that depend on both the quasiprimary operator $\mathcal{X}$ and the energy in the two states.
2.6 Contributions from $JJJJ$ in free massless boson theory
In free massless boson theory, there is the the current operator $J$, which is primary and has normalization $\alpha_{J}=1$ and conformal weights $(1,0)$. The OPE coefficient of CFTn primary operator $J_{j_{1}}J_{j_{2}}J_{j_{3}}J_{j_{4}}$ with $0\leq j_{1}<j_{2}<j_{3}<j_{4}\leq n-1$ is [42]
$d_{JJJJ}^{j_{1}j_{2}j_{3}j_{4}}=\frac{1}{(2n)^{4}}\bigg[\frac{1}{\big(\sin\frac{\pi j_{12}}{n}\sin\frac{\pi j_{34}}{n}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi j_{13}}{n}\sin\frac{\pi j_{24}}{n}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi j_{14}}{n}\sin\frac{\pi j_{23}}{n}\big)^{2}}\bigg].$ (2.38)
After summing over replica indices and organizing terms by their dependence on the expectation values, we obtain
$\mathcal{F}_{JJJJ}^{(p)}=\frac{1}{16}\big[(\langle J\rangle_{\rho}^{4}+\langle J\rangle_{\sigma}^{4})G_{6}(p)+(\langle J\rangle_{\rho}^{2}+\langle J\rangle_{\sigma}^{2})\langle J\rangle_{\rho}\langle J\rangle_{\sigma}G_{7}(p)+\langle J\rangle_{\rho}^{2}\langle J\rangle_{\sigma}^{2}G_{8}(p)\big]+O(n-1),$ (2.39)
where $G_{6}(p)$, $G_{7}(p)$, and $G_{8}(p)$ are defined in (A.18), (A.19), and (A), respectively. Using the analytical continuations (A.20), we obtain contributions from JJJJJJJJ
$\mathcal{F}_{JJJJ}^{(1/2)}=-\frac{1}{8192}(\langle J\rangle_{\rho}-\langle J\rangle_{\sigma})^{2}(19\langle J\rangle_{\rho}^{2}+26\langle J\rangle_{\rho}\langle J\rangle_{\sigma}+19\langle J\rangle_{\sigma}^{2}).$ (2.40)
The operator $\bar{J}$ has normalization $\alpha_{\bar{J}}=1$ and conformal weights $(0,1)$. Similarly, we obtain contributions from $\bar{J}\bar{J}\bar{J}\bar{J}$:
$\mathcal{F}_{\bar{J}\bar{J}\bar{J}\bar{J}}^{(1/2)}=-\frac{1}{8192}(\langle\bar{J}\rangle_{\rho}-\langle\bar{J}\rangle_{\sigma})^{2}(19\langle\bar{J}\rangle_{\rho}^{2}+26\langle\bar{J}\rangle_{\rho}\langle\bar{J}\rangle_{\sigma}+19\langle\bar{J}\rangle_{\sigma}^{2}).$ (2.41)
These quartic contributions provide higher-order corrections specific to the free boson theory and demonstrate in part how current operators contribute to subsystem fidelity.
2.7 Contributions from $\varepsilon\varepsilon\varepsilon\varepsilon$ in free massless fermion theory
In free massless fermion theory, the energy operator $\varepsilon$ is primary and has normalization $\alpha_{\varepsilon}=1$ and conformal weights $(h,\bar{h})=(\frac{1}{2},\frac{1}{2})$. From the correlation function on the plane
$\langle\varepsilon(z_{1},\bar{z}_{1})\varepsilon(z_{2},\bar{z}_{2})\varepsilon(z_{3},\bar{z}_{3})\varepsilon(z_{4},\bar{z}_{4})\rangle_{C}=\Big|\frac{1}{z_{12}z_{34}}-\frac{1}{z_{13}z_{24}}+\frac{1}{z_{14}z_{23}}\Big|^{2},$ (2.42)
we obtain the OPE coefficient of CFTn primary operator $\varepsilon_{j_{1}}\varepsilon_{j_{2}}\varepsilon_{j_{3}}\varepsilon_{j_{4}}$ with $0\leq j_{1}<j_{2}<j_{3}<j_{4}\leq n-1$
$d_{\varepsilon\varepsilon\varepsilon\varepsilon}^{j_{1}j_{2}j_{3}j_{4}}=\frac{1}{(2n)^{4}}\bigg(\frac{1}{\sin\frac{\pi j_{12}}{n}\sin\frac{\pi j_{34}}{n}}-\frac{1}{\sin\frac{\pi j_{13}}{n}\sin\frac{\pi j_{24}}{n}}+\frac{1}{\sin\frac{\pi j_{14}}{n}\sin\frac{\pi j_{23}}{n}}\bigg)^{2}.$ (2.43)
The summation over replica indices yields
$\mathcal{F}_{\varepsilon\varepsilon\varepsilon\varepsilon}^{(p)}=\frac{1}{16}\big[(\langle\varepsilon\rangle_{\rho}^{4}+\langle\varepsilon\rangle_{\sigma}^{4})G_{9}(p)+(\langle\varepsilon\rangle_{\rho}^{2}+\langle\varepsilon\rangle_{\sigma}^{2})\langle\varepsilon\rangle_{\rho}\langle\varepsilon\rangle_{\sigma}G_{10}(p)+\langle\varepsilon\rangle_{\rho}^{2}\langle\varepsilon\rangle_{\sigma}^{2}G_{11}(p)\big]+O(n-1),$ (2.44)
where $G_{9}(p)$, $G_{10}(p)$, and $G_{11}(p)$ are defined in (A.21), (A.22), and (A.23), respectively. Using the analytical continuations (A.24), we obtain contributions from $\varepsilon\varepsilon\varepsilon\varepsilon$
$\mathcal{F}_{\varepsilon\varepsilon\varepsilon\varepsilon}^{(1/2)}=-\frac{1}{8192}(\langle\varepsilon\rangle_{\rho}-\langle\varepsilon\rangle_{\sigma})^{2}(19\langle\varepsilon\rangle_{\rho}^{2}+26\langle\varepsilon\rangle_{\rho}\langle\varepsilon\rangle_{\sigma}+19\langle\varepsilon\rangle_{\sigma}^{2}).$ (2.45)
This result for the fermionic theory parallels the bosonic case in the previous subsection.
3 Free massless compact boson theory
The 2D free massless compact boson theory is a CFT with central charge $c=1$. We consider various low-lying primary states, including the ground state $|G\rangle$ with conformal weights $(0,0)$, vertex operator state $|\alpha,\bar{\alpha}\rangle=|V_{\alpha,\bar{\alpha}}\rangle$ with conformal weights $(\frac{\alpha^{2}}{2},\frac{\bar{\alpha}^{2}}{2})$, current state $|J\rangle$ with conformal weight $(1,0)$, anti-current state $|\bar{J}\rangle$ with conformal weight $(0,1)$, and state $|J\bar{J}\rangle$ with conformal weights $(1,1)$. Note that the ground state $|G\rangle=|0,0\rangle$. For current operators, we have expectation values:
$\langle J\rangle_{\alpha,\bar{\alpha}}=\frac{2\pi\mathrm{i}\alpha}{L},~~\langle\bar{J}\rangle_{\alpha,\bar{\alpha}}=-\frac{2\pi\mathrm{i}\bar{\alpha}}{L},$ (3.1)
$\langle J\rangle_{J}=\langle J\rangle_{\bar{J}}=\langle J\rangle_{J\bar{J}}=\langle\bar{J}\rangle_{J}=\langle\bar{J}\rangle_{\bar{J}}=\langle\bar{J}\rangle_{J\bar{J}}=0.$
Subsystem fidelities between these states have been obtained in [7, 9, 10]. We classify their short interval expansion calculations into three cases based on the expectation values in these states.
3.1 Case I
For case I, we have the following exact results
$F(\rho_{A,G},\rho_{A,J})=F(\rho_{A,G},\rho_{A,\bar{J}})=F(\rho_{A,J},\rho_{A,J\bar{J}})=F(\rho_{A,\bar{J}},\rho_{A,J\bar{J}})=\frac{\Gamma^{2}(\frac{3+\csc\frac{\pi x}{2}}{4})}{\Gamma^{2}(\frac{1+\csc\frac{\pi x}{2}}{4})}2\sin(\pi x),$
$F(\rho_{A,J},\rho_{A,\bar{J}})=F(\rho_{A,G},\rho_{A,J\bar{J}})=\frac{\Gamma^{4}(\frac{3+\csc\frac{\pi x}{2}}{4})}{\Gamma^{4}(\frac{1+\csc\frac{\pi x}{2}}{4})}4\sin^{2}(\pi x),$ (3.2)
where $x=\ell/L$ is the normalized interval length.
Using the universal contributions derived in the previous section, specifically equations (2.27) and (2.33), and the factorization formula
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{4}\mathcal{F}^{(1/2)}{TT}+\ell^{6}\mathcal{F}^{(1/2)}{TTT}+O(\ell^{8})\big]\big[1+\ell^{4}\mathcal{F}^{(1/2)}{\bar{T}\bar{T}}+\ell^{6}\mathcal{F}^{(1/2)}{\bar{T}\bar{T}\bar{T}}+O(\ell^{8})\big]$
[…]
we reproduce the short interval expansion of the above exact results. This agreement validates our approach and demonstrates that the dominant contributions in these cases come from the stress tensor and its descendants.
3.2 Case II
For case II, involving the ground state vertex operator states, we have the closed-form expression
$F(\rho_{A,\alpha,\bar{\alpha}},\rho_{A,\alpha^{\prime},\bar{\alpha}^{\prime}})=\Big(\cos\frac{\pi x}{2}\Big)^{\frac{(\alpha-\alpha^{\prime})^{2}+(\bar{\alpha}-\bar{\alpha}^{\prime})^{2}}{2}}.$ (3.4)
Using the contributions from current operators (2.27), (2.33), (2.40), and (2.41), and organizing them in a factorized form:
$F(\rho_{A},\sigma_{A})=\big[1+\ell^{2}\mathcal{F}^{(1/2)}{JJ}+\ell^{4}\big(\mathcal{F}^{(1/2)}{JJJJ}+\mathcal{F}^{(1/2)}{TJJ}+\mathcal{F}^{(1/2)}{TT}\big)+O(\ell^{6})\big]$
[…]
we reproduce the short interval expansion of the vertex operator fidelity. The factorization reflects the holomorphic-antiholomorphic separation in the free boson theory.
3.3 Case III
For case III, involving mixed states, we have the leading order results
$F(\rho_{A,J},\rho_{A,\alpha,\bar{\alpha}})=F(\rho_{A,\bar{J}},\rho_{A,\bar{\alpha},\alpha})=1-(\alpha^{2}+\bar{\alpha}^{2})\frac{\pi^{2}x^{2}}{16}+o(x^{2}),$
$F(\rho_{A,J\bar{J}},\rho_{A,\alpha,\bar{\alpha}})=1-(\alpha^{2}+\bar{\alpha}^{2})\frac{\pi^{2}x^{2}}{16}+o(x^{2}),$ (3.6)
Using the expansion (3.2), we obtain results with higher-order terms
$F(\rho_{A,J},\rho_{A,\alpha,\bar{\alpha}})=F(\rho_{A,\bar{J}},\rho_{A,\bar{\alpha},\alpha})=1-\frac{(\alpha^{2}+\bar{\alpha}^{2})\pi^{2}x^{2}}{16}$
[…]
$F(\rho_{A,J\bar{J}},\rho_{A,\alpha,\bar{\alpha}})=1-\frac{(\alpha^{2}+\bar{\alpha}^{2})\pi^{2}x^{2}}{16}+\frac{[3(\alpha^{2}+\bar{\alpha}^{2})^{2}+44(\alpha^{2}+\bar{\alpha}^{2})-288]\pi^{4}x^{4}}{1536}+O(x^{6}).$
No higher-order analytical results exist in the literature for these cases, so we compare our predictions with numerical results in the spin-1/2 XX chain. In the XX chain, the above CFT results are expected to apply in the limit $1\ll\ell\ll L$. To match CFT and spin chain results, we require high-efficiency and high-precision numerical evaluation of subsystem fidelity in the spin chain. We calculate subsystem fidelity in the XX chain using the truncation method with truncation number $t=8$ from [43] and show results in Figure 1. We define the subtracted fidelity
$F_{(2)}\equiv 1-\frac{(\alpha^{2}+\bar{\alpha}^{2})\pi^{2}x^{2}}{16},$ (3.8)
which is removed from both CFT and spin chain results for clearer comparison. The figure uses $(\alpha,\bar{\alpha})=(1,0)$, showing good agreement between CFT predictions and spin chain results. We also calculated cases with other values of $(\alpha,\bar{\alpha})$, though the agreement is not always as good, possibly due to finite-size effects or limitations of the numerical method.
Figure 1: Comparison between CFT predictions (solid red lines) and $X$ spin chain results (symbols) for the subtracted fidelity $F_{(2)}-F$. The good agreement validates the short-interval expansion approach.
4 Free massless fermion theory
The 2D free massless fermion theory is a CFT with central charge $c=\frac{1}{2}$. We consider various low-lying primary states, including the ground state $|G\rangle$ with conformal weights $(h,\bar{h})=(0,0)$; the states $|\sigma\rangle$ and $|\mu\rangle$, created by the spin operator $\sigma$ and disorder operator $\mu$, respectively, both with conformal weights $(\frac{1}{16},\frac{1}{16})$; the state $|\psi\rangle$ with conformal weight $(\frac{1}{2},0)$; the state $|\bar{\psi}\rangle$ with conformal weight $(0,\frac{1}{2})$; and the state $|\epsilon\rangle$, created by the energy operator $\epsilon$, with conformal weights $(\frac{1}{2},\frac{1}{2})$. Expectation values for the energy operator are
$\langle\varepsilon\rangle_{G}=\langle\varepsilon\rangle_{\psi}=\langle\varepsilon\rangle_{\bar{\psi}}=\langle\varepsilon\rangle_{\varepsilon}=0,~~\langle\varepsilon\rangle_{\sigma}=-\langle\varepsilon\rangle_{\mu}=\frac{\pi}{L},$ (4.1)
We classify related subsystem fidelities between these states from [9, 10] into three cases.
4.1 Case I
For case I, we have the exact results
$F(\rho_{A,G},\rho_{A,\psi})=F(\rho_{A,G},\rho_{A,\bar{\psi}})=F(\rho_{A,\psi},\rho_{A,\varepsilon})=F(\rho_{A,\bar{\psi}},\rho_{A,\varepsilon})=\frac{\Gamma(\frac{3+\csc\frac{\pi x}{2}}{4})}{\Gamma(\frac{1+\csc\frac{\pi x}{2}}{4})}\sqrt{2\sin(\pi x)},$
$F(\rho_{A,\psi},\rho_{A,\bar{\psi}})=F(\rho_{A,G},\rho_{A,\varepsilon})=\frac{\Gamma^{2}(\frac{3+\csc\frac{\pi x}{2}}{4})}{\Gamma^{2}(\frac{1+\csc\frac{\pi x}{2}}{4})}2\sin(\pi x).$ (4.2)
Using the same factorization formula (3.1) as in the bosonic case, we reproduce the short interval expansion of these results.
4.2 Case II
For case II, involving the ground state and the spin and disorder states, we have
$F(\rho_{A,G},\rho_{A,\sigma})=F(\rho_{A,G},\rho_{A,\mu})=\Big(\cos\frac{\pi x}{2}\Big)^{\frac{1}{8}},$
$F(\rho_{A,\sigma},\rho_{A,\mu})=\Big(\cos\frac{\pi x}{2}\Big)^{\frac{1}{2}}.$ (4.3)
Using the expansion organized around the energy operator $\varepsilon$
$F(\rho_{A},\sigma_{A})=1+\ell^{2}\mathcal{F}^{(1/2)}{\varepsilon\varepsilon}+\ell^{4}\big(\mathcal{F}^{(1/2)}{\varepsilon\varepsilon\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{T\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{\bar{T}\varepsilon\varepsilon}+\mathcal{F}^{(1/2)}{TT}+\mathcal{F}^{(1/2)}{\bar{T}\bar{T}}\big)+O(\ell^{6}),$ (4.4)
we reproduce the short interval expansion of the above results.
4.3 Case III
For case III, involving mixed states between spin and disorder operators and other primary states, we have the leading order results
$F(\rho_{A,\sigma},\rho_{A,\psi})=F(\rho_{A,\sigma},\rho_{A,\bar{\psi}})=F(\rho_{A,\mu},\rho_{A,\psi})=F(\rho_{A,\mu},\rho_{A,\bar{\psi}})=1-\frac{\pi^{2}x^{2}}{64}+o(x^{2}),$
$F(\rho_{A,\sigma},\rho_{A,\varepsilon})=F(\rho_{A,\mu},\rho_{A,\varepsilon})=1-\frac{\pi^{2}x^{2}}{64}+o(x^{2}),$ (4.5)
Using expansion (4.4), we obtain higher-order corrections
$F(\rho_{A,\sigma},\rho_{A,\psi})=F(\rho_{A,\sigma},\rho_{A,\bar{\psi}})=F(\rho_{A,\mu},\rho_{A,\psi})=F(\rho_{A,\mu},\rho_{A,\bar{\psi}})=1-\frac{\pi^{2}x^{2}}{64}-\frac{1069\pi^{4}x^{4}}{24576}+O(x^{6}),$
$F(\rho_{A,\sigma},\rho_{A,\varepsilon})=F(\rho_{A,\mu},\rho_{A,\varepsilon})=1-\frac{\pi^{2}x^{2}}{64}-\frac{2125\pi^{4}x^{4}}{24576}+O(x^{6}).$ (4.6)
We compare these predictions with numerical results in the critical Ising chain, obtained using the truncation method from [43] with truncation number $t=8$, as shown in Figure 2. We define the subtracted fidelity
$F_{(2)}\equiv 1-\frac{\pi^{2}x^{2}}{64},$ (4.7)
which is removed from both CFT and spin chain results. The agreement between our analytical predictions and numerical simulations provides strong validation of our approach.
Figure 2: Comparison between CFT predictions (solid red lines) and critical Ising chain results (symbols) for the subtracted fidelity $F_{(2)}-F$. The agreement demonstrates the applicability of our method to fermionic systems.
5 2D holographic CFTs
The 2D holographic CFTs have large central charge $c=\frac{3R}{2G}$ [35], with $G$ being the Newton constant and $R$ being the AdS radius, and sparse low-lying spectrum [44]. We focus only on contributions from the vacuum conformal family, where holomorphic and anti-holomorphic sectors factorize
$F(\rho_{A},\sigma_{A})=F_{\rm{holo}}(\rho_{A},\sigma_{A})F_{\rm{anti-holo}}(\rho_{A},\sigma_{A}).$ (5.1)
This section shows only contributions from the holomorphic sector only, as the anti-holomorphic sector follows by analogy. Part of the results in this section has been reported in [41]. Here, we present higher-order results and additional calculation details.
Consider two primary states $|\phi\rangle$ and $|\psi\rangle$ with large conformal weights $h_{\phi}=c\epsilon_{\phi}$ and $h_{\psi}=c\epsilon_{\psi}$, respectively. Here $\epsilon_{\phi}\sim O(c^{0})$, $\epsilon_{\psi}\sim O(c^{0})$, $\epsilon_{\phi}-\epsilon_{\psi}\sim O(c^{0})$. Using our universal formulas (2.27) and (2.33) with $\mathcal{X}=T$ and keeping only the contributions from the holomorphic sector of the identity conformal family, we obtain the fidelity expansion
$F(\rho_{A,\phi},\rho_{A,\psi})=1-\frac{3\pi^{4}c\ell^{4}(\epsilon_{\phi}-\epsilon_{\psi})^{2}}{32L^{4}}+\frac{\pi^{6}c\ell^{6}(\epsilon_{\phi}-\epsilon_{\psi})^{2}(12\epsilon_{\psi}+12\epsilon_{\phi}-1)}{64L^{6}}+O(\ell^{8}).$ (5.2)
This result shows that for two heavy states with large different energies in holographic CFTs, the fidelity deviation from unity is of order $O(c)$, reflecting the semiclassical nature of these states in the holographic dual.
Now consider primary states $|\phi\rangle$ and $|\psi\rangle$ with the same leading order conformal weights $h_{\phi}=c\epsilon_{\phi}+\delta_{\phi}$ and $h_{\phi}=c\epsilon_{\phi}+\delta_{\psi}$, where $\delta_{\phi}\sim O(c^{0})$, $\delta_{\psi}\sim O(c^{0})$, and $\delta_{\phi}-\delta_{\psi}\sim O(c^{0})$, represent small excitations above a heavy background. In this case, we obtain
$F(\rho_{A,\phi},\rho_{A,\psi})=1-\frac{3\pi^{4}\ell^{4}(\delta_{\phi}-\delta_{\psi})^{2}}{32cL^{4}}+\frac{\pi^{6}\ell^{6}(\delta_{\phi}-\delta_{\psi})^{2}[c(24\epsilon_{\phi}-1)+12(\delta_{\psi}+\delta_{\phi})]}{64c^{2}L^{6}}+O(\ell^{8}).$ (5.3)
At leading order, the two black hole microstates have the same energy and are classically indistinguishable; however, perturbative $1/c$ quantum corrections lift this degeneracy and allow them to be distinguished.
Finally, we consider a primary state $|\phi\rangle$ with conformal weight $h_{\phi}=c\epsilon_{\phi}$ and a thermal state $\rho_{\beta}=\frac{\mathrm{e}^{-\beta H}}{Z(\beta)}$ with inverse temperature $\beta$ and $Z(\beta)=\textrm{tr}(\mathrm{e}^{-\beta H})$. In the case where [45, 46]
$\beta=\frac{L}{\sqrt{24\epsilon_{\phi}-1}},$ (5.4)
we have $\langle T\rangle_{\phi}=\langle T\rangle_{\beta}$, meaning the states have matching one-point functions of the stress tensor. In this finely tuned situation, we use the our universal formula (2.27) with $\mathcal{X}=\mathcal{A}$ and obtain
$F(\rho_{A,\phi},\rho_{A,\beta})=1-\frac{7\pi^{8}c\ell^{8}(22\epsilon_{\phi}-1)^{2}\epsilon_{\phi}^{2}}{512(5c+22)L^{8}}+O(\ell^{10}).$ (5.5)
This work provides direct evidence for the perturbative distinguishability between black hole microstates and the canonical ensemble thermal state, and this contrasts with the indirect evidence from earlier studies [47, 28, 30, 31, 32]
6 Conclusion
In this work, we have investigated subsystem fidelity in 2D CFTs using the short-interval expansion derived from the OPE of twist operators. Our approach provides a comprehensive framework for computing fidelity between reduced density matrices of various states in 2D CFTs. We derived explicit expressions for universal contributions from general families of quasiprimary operators, including $\mathcal{X}\mathcal{X}$, $\mathcal{X}\mathcal{X}\mathcal{X}$, and $T\mathcal{X}\mathcal{X}$, in arbitrary 2D CFTs, capturing both leading and subleading behavior in the short-interval limit. We computed specific contributions from particular quasiprimary operators in free massless boson and fermion theories, demonstrating the method’s applicability to concrete examples. Our results show excellent agreement with known analytical expressions and numerical simulations in integrable lattice models, validating the proposed approach. We extended the method to 2D holographic CFTs, analyzing the perturbative distinguishability of black hole microstates through the AdS/CFT correspondence.
The short-interval expansion method developed here offers several advantages: it provides a systematic way to compute fidelity order by order in the interval length, reveals the operator content responsible for state distinguishability. The factorization between holomorphic and antiholomorphic sectors in many cases further simplifies computations and provides physical insight. However, several limitations remain. The expansion is inherently perturbative, and its convergence properties at larger intervals require further investigation. Discrepancies observed in certain cases involving vertex operator states may stem from numerical precision limitations, finite-size effects, or OPE truncation subtleties. Additionally, our holographic treatment relies on the vacuum dominance approximation, which may not capture all significant contributions in finite central charge regimes, while the analytical continuation procedure becomes increasingly complex when higher-order quasiprimary operators are included.
Looking forward, several promising research directions emerge. Extending the short-interval expansion to higher orders by incorporating additional quasiprimary operators would improve accuracy over wider interval ranges. Applying this method to other CFT models, such as supersymmetric theories, non-unitary CFTs, or theories with extended symmetry algebras, could reveal new universal features of subsystem fidelity. Refining the approach for holographic settings may lead to deeper understanding of black hole microstate distinguishability and the information paradox. Furthermore, generalizing the formalism to disjoint intervals, higher-dimensional CFTs, or out-of-equilibrium states would potentially broaden its applicability across quantum information and gravitational physics.
Acknowledgements
J.Z. is grateful to the organizers and participants of the Annual Academic Conference of the Gravitation and Relativistic Astrophysics Division of the Chinese Physical Society (Kunming, April 2025), the Symposium on Gravity and Cosmology (Xi’an, June 2025), and the Symposium on Strings, Gravity, and Gravitational Waves (Rizhao, August 2025), where preliminary results of this work were presented and benefitted from valuable discussions. J.Z. also wishes to thank the Institute of Fundamental Physics and Quantum Technology at Ningbo University for its hospitality during a productive visit where part of this research was presented and completed. In addition, J.Z. thanks Bin Chen, Wu-zhong Guo, Song He, Yan Liu, and Jie-qiang Wu for insightful discussions and constructive comments.
This work is supported by the National Natural Science Foundation of China (Grant No. 12205217) and the Tianjin University Self-Innovation Fund Extreme Basic Research Project (Grant No. 2025XJ21-0007).
Appendix A Analytical continuation
In the appendix, we collect the analytical continuation for several summation formulas.
We begin by defining
$G_{1}(p,\Delta)=\sum_{a_{1}<a_{2}}\frac{1}{\big|\sin\frac{\pi a_{12}}{2p}\big|^{2\Delta}},$ (A.1)
where $a_{12}\equiv a_{1}-a_{2}$ and $a_{1},a_{2}$ take even integer values in the set $\{0,2,\cdots,2p-2\}$. Our goal is to obtain $G_{1}(\frac{1}{2},\Delta)$ for general $\Delta$. First, we compute
$G_{1}(p,1)=\frac{p(p^{2}-1)}{6},$ (A.2)
which gives
$G_{1}\Big(\frac{1}{2},1\Big)=-\frac{1}{16}.$ (A.3)
Proceeding similarly, we obtain the following values
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|c| }\hline\cr\Delta&1&2&3&4&5&6&7&8&9&\cdots\\ \hline\cr G_{1}(\frac{1}{2},\Delta)&-\frac{1}{16}&-\frac{3}{64}&-\frac{5}{128}&-\frac{35}{1024}&-\frac{63}{2048}&-\frac{231}{8192}&-\frac{429}{16384}&-\frac{6435}{262144}&-\frac{12155}{524288}$$\cdots\\ \hline\cr\end{array}$ (A.4)
From these values, we extract the general formula
$G_{1}\Big(\frac{1}{2},\Delta\Big)=-\frac{\Gamma(\Delta+\frac{1}{2})}{8\sqrt{\pi}\Gamma(\Delta+1)},$ (A.5)
which can be verified for further values of $\Delta$.
Next, we define
$G_{2}(p,\Delta)=\sum_{a_{1}<a_{2}<a_{3}}\frac{1}{\big|\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi a_{13}}{2p}\sin\frac{\pi a_{23}}{2p}\big|^{\Delta}},$ (A.6)
Using the same method as above, we compute
$\Delta$ 2 4 6 8 10 12 14 16
$G_{2}(\frac{1}{2},\Delta)$ $\frac{3}{128}$ $\frac{245}{8192}$ $\frac{12705}{262144}$ $\frac{2927925}{33554432}$ $\frac{44757141}{268435456}$ $\frac{22748036311}{68719476736}$ $\frac{742182172875}{1099511627776}$ $\frac{790292983267125}{562949953421312}$
18 20 22 24 $\cdots$
$\frac{13349421931027875}{4503599627370496}$ $\frac{1825735334414506515}{288230376151711744}$ $\frac{504282766251826384095}{36893488147419103232}$ $\frac{140415205544153506193175}{4722366482869645213696}$ $\cdots$
From these results, we extract the general expression
$G_{2}\Big(\frac{1}{2},\Delta\Big)=\frac{2^{\Delta-4}\Gamma(\Delta+\frac{1}{2})^{2}}{\pi\Gamma(\frac{\Delta}{2}+1)\Gamma(\frac{3\Delta}{2}+1)}.$ (A.8)
We now introduce
$G_{3}(p,\Delta)=\sum_{a_{1}<a_{2}<a_{3}}\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi a_{13}}{2p}\big)^{2}\big|\sin\frac{\pi a_{23}}{2p}\big|^{2\Delta_{\mathcal{X}}-2}},$ (A.9)
and obtain the values
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|c| }\hline\cr\Delta&1&2&3&4&5&6&7&8&9&\cdots\\ \hline\cr G_{3}(\frac{1}{2},\Delta)&\frac{1}{16}&\frac{9}{128}&\frac{35}{512}&\frac{133}{2048}&\frac{63}{1024}&\frac{957}{16384}&\frac{7293}{131072}&\frac{27885}{524288}&\frac{26741}{524288}$$\cdots\\ \hline\cr\end{array}$ (A.10)
from which we derive
$G_{3}\Big(\frac{1}{2},\Delta\Big)=\frac{(5\Delta-1)\Gamma(\Delta+\frac{1}{2})}{16\sqrt{\pi}\Gamma(\Delta+2)}.$ (A.11)
We further define
$G_{4}(p,\Delta)=\sum_{a_{1}<a_{2},b}\frac{1}{\big[\sin\frac{\pi(a_{1}-b)}{2p}\sin\frac{\pi(a_{2}-b)}{2p}\big]^{2}\big|\sin\frac{\pi a_{12}}{2p}\big|^{2\Delta_{\mathcal{X}}-2}},$ (A.12)
and compute
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|c| }\hline\cr\Delta&1&2&3&4&5&6&7&8&9&\cdots\\ \hline\cr G_{4}(\frac{1}{2},\Delta)&-\frac{1}{32}&-\frac{3}{128}&-\frac{5}{256}&-\frac{35}{2048}&-\frac{63}{4096}&-\frac{231}{16384}&-\frac{429}{32768}&-\frac{6435}{524288}&-\frac{12155}{1048576}$$\cdots\\ \hline\cr\end{array}$ (A.13)
leading to
$G_{4}\Big(\frac{1}{2},\Delta\Big)=-\frac{\Gamma(\Delta+\frac{1}{2})}{16\sqrt{\pi}\Gamma(\Delta+1)}.$ (A.14)
Next, we consider
$G_{5}(p,\Delta)=\sum_{a_{1}\neq a_{2},b}\frac{1}{\big[\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi(a_{1}-b)}{2p}\big]^{2}\big|\sin\frac{\pi(a_{2}-b)}{2p}\big|^{2\Delta-2}},$ (A.15)
and find
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|c| }\hline\cr\Delta&1&2&3&4&5&6&7&8&9&\cdots\\ \hline\cr G_{5}(\frac{1}{2},\Delta)&-\frac{1}{32}&-\frac{3}{64}&-\frac{25}{512}&-\frac{49}{1024}&-\frac{189}{4096}&-\frac{363}{8192}&-\frac{5577}{131072}&-\frac{10725}{262144}&-\frac{41327}{1048576}$$\cdots\\ \hline\cr\end{array}$ (A.16)
from which we extract
$G_{5}\Big(\frac{1}{2},\Delta\Big)=-\frac{(2\Delta-1)\Gamma(\Delta+\frac{1}{2})}{8\sqrt{\pi}\Gamma(\Delta+2)}.$ (A.17)
We also introduce the definitions
$\hskip-28.45274ptG_{6}(p)=\sum_{a_{1}<a_{2}<a_{3}<a_{4}}\bigg[\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi a_{34}}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi a_{13}}{2p}\sin\frac{\pi a_{24}}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi a_{14}}{2p}\sin\frac{\pi a_{23}}{2p}\big)^{2}}\bigg],$ (A.18)
$\hskip-28.45274ptG_{7}(p)=\sum_{a_{1}<a_{2}<a_{3},b}\bigg[\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi(a_{3}-b)}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi a_{13}}{2p}\sin\frac{\pi(a_{2}-b)}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b)}{2p}\sin\frac{\pi a_{23}}{2p}\big)^{2}}\bigg],$ (A.19)
$\hskip-28.45274ptG_{8}(p)=\sum_{a_{1}<a_{2},b_{1}<b_{2}}\bigg[\frac{1}{\big(\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi b_{12}}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{1})}{2p}\sin\frac{\pi(a_{2}-b_{2})}{2p}\big)^{2}}+\frac{1}{\big(\sin\frac{\pi(a_{1}-b_{2})}{2p}\sin\frac{\pi(a_{2}-b_{1})}{2p}\big)^{2}}\bigg],\$
with the corresponding analytical continuations
$G_{6}\Big(\frac{1}{2}\Big)=-\frac{19}{512},~~G_{7}\Big(\frac{1}{2}\Big)=\frac{3}{128},~~G_{8}\Big(\frac{1}{2}\Big)=\frac{7}{256}.$ (A.20)
Finally, we define
$\hskip-45.5244ptG_{9}(p)=\sum_{a_{1}<a_{2}<a_{3}<a_{4}}\bigg(\frac{1}{\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi a_{34}}{2p}}-\frac{1}{\sin\frac{\pi a_{13}}{2p}\sin\frac{\pi a_{24}}{2p}}+\frac{1}{\sin\frac{\pi a_{14}}{2p}\sin\frac{\pi a_{23}}{2p}}\bigg)^{2},$ (A.21)
$\hskip-45.5244ptG_{10}(p)=\sum_{a_{1}<a_{2}<a_{3},b}\bigg(\frac{1}{\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi(a_{3}-b)}{2p}}-\frac{1}{\sin\frac{\pi a_{13}}{2p}\sin\frac{\pi(a_{2}-b)}{2p}}+\frac{1}{\sin\frac{\pi(a_{1}-b)}{2p}\sin\frac{\pi a_{23}}{2p}}\bigg)^{2},$ (A.22)
$\hskip-45.5244ptG_{11}(p)=\sum_{a_{1}<a_{2},b_{1}<b_{2}}\bigg(\frac{1}{\sin\frac{\pi a_{12}}{2p}\sin\frac{\pi b_{12}}{2p}}-\frac{1}{\sin\frac{\pi(a_{1}-b_{1})}{2p}\sin\frac{\pi(a_{2}-b_{2})}{2p}}+\frac{1}{\sin\frac{\pi(a_{1}-b_{2})}{2p}\sin\frac{\pi(a_{2}-b_{1})}{2p}}\bigg)^{2},$ (A.23)
and obtain the analytical continuations
$G_{9}\Big(\frac{1}{2}\Big)=-\frac{19}{512},~~G_{10}\Big(\frac{1}{2}\Big)=\frac{3}{128},~~G_{11}\Big(\frac{1}{2}\Big)=\frac{7}{256}.$ (A.24)